Chemist prepares dilute aqueous solution of NH4Cl. This was achieved by diluting 25mL of a concentrated solution to a final volume of 3.5L. The concentration of NH4Cl in the dilute solution was 0.0659M. The original concentrated solution was prepared by dissolving 74.027g of the solid chloride in water and making up the solution in a volumetric flask. Calculate the volume of the flask used to make up the concentrated solution.
1
Expert's answer
2018-01-16T02:02:56-0500
Solution:
Let’s first calculate amount of NH4Cl in the dilute solution.
n = C*V = 0.23065 moles.
This amount is equal to amount of NH4Cl in the concentrated solution with volume of 25mL.
So, let’s calculate the concentration of NH4Cl in the concentrated solution.
C = n/V = 0.23065/0.025 = 9.226 M.
When preparing the solution, we took 74.027g, which is 74.027/53.5 = 1.38 moles of solid ammonium chloride. Dissolving this amount of chloride in water, we get a solution with concentration of NH4Cl equal to C. Then, when putting that into a volumetric flask, the concentration of NH4Cl in the solution doesn’t change, it’s still equal to C.
Hence, we can find volume of the flask if we divide the amount of NH4Cl (1.38 moles) by its concentration (9.226 M).
V = 1.38/9.226=0.15 L.
Answer: 150mL